Emily is installing artificial grass in her garden. The garden is 15 feet long and 30 feet wide. It took her 4 hours to install the grass.
What is the average speed at which she installed the grass?

Answer by TimothyLamb(4379)

a = 15*30 = 450 sq.ft
---
r = a / t
---
r = 450 / 4
---
r = 112.5 sq.ft/hr
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php